OpenStudy (anonymous):

Two forces with magnitudes of 150 and 75 pounds act on an object at angles of 30° and 150° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

you change it to x and y coordinates for each forces.

OpenStudy (anonymous):

x direction of 150 pound force= (150/2.2)g cos30 y direction of 150 pound force= (150/2.2)gsin30

OpenStudy (anonymous):

x direction of 75 pond force= (75/2.2)g cos150 y direction of 75 pound force= (75/2.2)g sin150. Add the x directions and y directions together. Force= sqrt(x force^2+ydorce^2) @ angle = arctan(y force/ x force) , where x force is resultant force in x axis, y is resultant force on the y axis.

OpenStudy (anonymous):

I dont quite understand what you mean by 150 and 75 pounds of force however.:(

OpenStudy (anonymous):

where do you get the 2.2 from though :/

OpenStudy (anonymous):

converting from pounds to kg. . . you divide pounds by 2.2 = kg.

OpenStudy (anonymous):

f= mg, (usually vertical force downward due to gravity) . but as i mentioned previously, I DONT UNDERSTAND why you have pounds not newtons ...

OpenStudy (anonymous):

idk it's what it gave me in the question... but so i add the sin and cos of 30 with the 150? and then i get 205

OpenStudy (anonymous):

nononono. you add x forces of 150 pounds and 75 pounds, and you add y forces of 150 pounds and 75 pounds.

OpenStudy (anonymous):

Then this will result x force and y force of the resultant. and the resultant force's magnitude is equivalent to getting distance of the resultant x resultant force and y resultant force == sqrt ( x resultant force^2+ y resultant force^2).

OpenStudy (anonymous):

angle of the resultant force is arctan(y resultant / x resultant force) .

OpenStudy (anonymous):

sqrt(112^2+65^2) ?

OpenStudy (anonymous):

= 129?

OpenStudy (anonymous):

and would the angle be 19?

OpenStudy (anonymous):

magnitude is correct.

OpenStudy (anonymous):

x resultant force is 65

OpenStudy (anonymous):

so then 30 as direction?

OpenStudy (anonymous):

so it should be arctan(112/65)

OpenStudy (anonymous):

recall: angle= arctan (y resultant force/ x resultant force)= arctan(112/65)

OpenStudy (anonymous):

60?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay!! thank you so much